### v2

# Ecology, Spatial Structure, and Selection Pressure Induce Strong Signatures in Phylogenetic Structure

Abstract
Evolutionary dynamics are shaped by a variety of fundamental, generic drivers, including spatial structure, ecology, and selection pressure. These drivers impact the trajectory of evolution, and have been hypothesized to influence phylogenetic structure. For instance, they can help explain natural history, steer behavior of contemporary evolving populations, and influence efficacy of application-oriented evolutionary optimization. Likewise, in inquiry-oriented artificial life systems, these drivers constitute key building blocks for open-ended evolution. Here, we set out to assess (1) if spatial structure, ecology, and selection pressure leave detectable signatures in phylogenetic structure, (2) the extent, in particular, to which ecology can be detected and discerned in the presence of spatial structure, and (3) the extent to which these phylogenetic signatures generalize across evolutionary systems. To this end, we analyze phylogenies generated by manipulating spatial structure, ecology, and selection pressure within three computational models of varied scope and sophistication. We find that selection pressure, spatial structure, and ecology have characteristic effects on phylogenetic metrics, although these effects are complex and not always intuitive. Signatures have some consistency across systems when using equivalent taxonomic unit definitions (e.g., individual, genotype, species). Further, we find that sufficiently strong ecology can be detected in the presence of spatial structure. We also find that, while low-resolution phylogenetic reconstructions can bias some phylogenetic metrics, high-resolution reconstructions recapitulate them faithfully. Although our results suggest potential for evolutionary inference of spatial structure, ecology, and selection pressure through phylogenetic analysis, further methods development is needed to distinguish these drivers’ phylometric signatures from each other and to appropriately normalize phylogenetic metrics. With such work, phylogenetic analysis could provide a versatile toolkit to study large-scale evolving populations.
Introduction
Within evolutionary biology, the structure of ancestry relationships among organisms within an evolving system is described as their “phylogeny.” Phylogenies detail the sequence of historical lineage-branching events that gave rise to contemporary populations. Most obviously, the phylogenetic record can reveal the ordering and timing of particular contingencies (e.g., trait originations, population separations) that gave rise to current populations (pagel1999inferring; arbogast2002estimating). For instance, phylogenies reconstructed from fungal gene sequences suggest frequent switching of nutritional sources and that spore-dispersal mechanisms have multiple independent origins (james2006reconstructing).
However, phylogenetic analyses can also test more general hypotheses about underlying evolutionary dynamics. Rather than seeking to detect specific events, researchers using phylogenies for this purpose instead measure structural patterns that may be indicative of different evolutionary regimes. For instance, analysis of the counts of coexisting lineages over time has been used to detect density-dependent changes in speciation rate (rabosky2008density), which can be interpreted as an indication of transition away from niche expansion due to lessening availability of ecological opportunity.
Indeed, methods to infer underlying evolutionary dynamics from phylogenetic structure are increasingly harnessed for tangible, impactful applications. In evolutionary oncology, for example, phylogenetic analysis has been used to quantify patterns of tumor evolution (scottInferringTumorProliferative2020; lewinsohnStatedependentEvolutionaryModels2023). Likewise, pathogen phylogenies have been studied to identify dynamics of epidemiological spread of Ebola and HIV (giardina2017inference; saulnier2017inferring; voznica2022deep). Phylogenetic information also has a long history of helping identify policy priorities for ecological conservation (forestPreservingEvolutionaryPotential2007).
In computational evolutionary systems, phylogenetic structure metrics are a powerful tool for summarizing evolutionary history (dolson2020interpreting). These analyses can help digital systems further our knowledge of biology, but they are also directly valuable for learning about digital systems themselves. Notably, within the realm of artificial life research, phylogeny-based metrics have been proposed to identify hallmarks of open-ended evolution (dolson2019modes). They have also been used to understand evolutionary trajectories that lead to outcomes of interest (lenskiEvolutionaryOriginComplex2003; lalejiniEvolutionaryOriginsPhenotypic2016; johnsonEndosymbiosisBustInfluence2022a). For evolution-inspired optimization algorithms, phylogeny-based metrics have been used to predict which runs of evolutionary computation will be successful (hernandezWhatCanPhylogenetic2022a; shahbandeganUntanglingPhylogeneticDiversity2022a; ferguson2023potentiating) and even proactively guide artificial selection algorithms (lalejini2024phylogeny; burke2003increased).
Here, we assess the potential of phylogenetic analysis as a means to study fundamental, universal evolutionary dynamics — ecology, selection pressure, and spatial structure. Specifically, we consider the following questions,
1.
Do these dynamics leave detectable signatures in phylogenetic structure?
2.
If so, to what extent are these dynamics’ signatures discernable from one another?
3.
To what extent do the structure of these dynamics’ signatures generalize across evolutionary systems?
We use in silico experiments to investigate these questions, and test across two levels of taxonomic abstraction. To quantify phylogenetic structure at organism-level abstraction — meaning that phylogeny is represented at the granularity of individual parent-child relationships — we use both a minimalistic, forward-time simulation and a rich artificial life system (ofria2004avida). In contrast, at species-level abstraction (i.e. species trees), we used a population-level simulation designed to study origins of biodiversity (hagen2021gen3sis).
A major strength of simulation experiments for this purpose is that they provide exact ground-truth phylogenies due to their perfectly observable evolutionary history. However, for in vivo systems perfect phylogeny tracking is not typically possible. Although new techniques are making increasingly high-fidelity lineage tracking possible in some experiments (nozoe2017inferring; woodworth2017building), phylogenies are still usually estimated from naturally-occuring biosequence data. Moreover, as digital evolution systems scale, issues of data loss and centralization overhead make perfect tracking at best inefficient and at worst untenable. Thus, some systems will likely need to adopt a decentralized, reconstruction-based approach similar to biological data (moreno2024analysis), which can be achieved through the recently-developed “hereditary stratigraphy” methodology (moreno2022hstrat).
Effective use of phylogenetic metrics in scenarios calculated from estimated phylogenies (as opposed to exact phylogenies) requires consideration of potential confounding effects from inaccuracies introduced by reconstruction. However, it is in precisely such scenarios where the new capabilities to characterize evolutionary dynamics could have the greatest impact; large-scale systems can produce an intractable quantity of data, making phylometrics valuable as summary statistics of the evolutionary process (dolson2020interpreting). In particular, methods are needed to handle large-scale artificial life systems where complete, perfect visibility is not feasible and evolution operates according to implicit, contextually dependent fitness dynamics (moreno2022exploring; kojima2023implementation). Thus, an additional goal of this paper is to quantify the magnitude and character of bias introduced by reconstruction error from hereditary stratigraphy. To this end, we calculate phylometric metrics across a range of reconstruction accuracy levels, and report the level of accuracy necessary to attain metric readings statistically indistinguishable from ground truth.
This work builds on a series of recent studies developing, testing, and applying hereditary stratigraphy methodology since its introduction in moreno2022hereditary. We utilized the open-source hstrat software library’s hereditary stratigraphy algorithm implementations, which are publicly available via the Python Packaging Index (moreno2022hstrat). In perfect-tracking experiments, we collected ground-truth baseline phylogenies using the Phylotrack library (dolson2024phylotrack). moreno2024guide characterized trade-offs between memory use, inference precision, and inference accuracy across hereditary stratigraphy configurations, providing a foundation for best practices in applying the methodology. The present paper focuses on establishing foundations for using herediatry stratigraphy to infer evolutionary dynamics from phylogenetic history. Specifically, the work extends a conference paper moreno2023toward, by adding (1) replications of experiments in full-fledged evolution simulation frameworks (i.e., Avida and Gen3sis, introduced in methods) and (2) refining the set of phylogeny metrics employed. Findings from present work have facilitated application of hereditary stratigraphy to characterize dynamics in evolution simulations run on the 850,000-core Cerebras Wafer-Scale Engine (moreno2024trackable), which required engineering a simpler and more efficient algorithmic basis for hereditary stratigraphy (moreno2024structured). Although present work considers non-hybridizing phylogenies (i.e., asexual ancestry trees and species trees), methods applying hereditary stratigraphy to sexual populations have been proposed in moreno2024methods.
The utility of a model system for experimental evolution hinges on sufficient ability to observe and interpret underlying evolutionary dynamics. Benchtop and field-based biological models continue to benefit from ongoing methodological advances that have profoundly increased visibility into genetic, phenotypic, and phylogenetic state (woodworth2017building; blomberg2011measuring; schneider2019past). Simulation systems, in contrast, have traditionally enjoyed perfect, complete observability (hindre2012new). However, ongoing advances in parallel and distributed computing hardware have begun to strain analytical omnipotence, as increasingly vast throughput introduces challenges centralizing, storing, and analyzing data (klasky2021data). These trends, although unfolding opposite one another, exhibit intriguing convergence: computational and biological domains will provide data that is multimodal and high-resolution but also incomplete and imperfect. Such convergence establishes great potential for productive interdisciplinary exchange.
Methods
Model Systems
To assess the generality of evolutionary dynamics’ effects on phylogenetic structure, we replicated experiments across models differing in subject and approach. We used three systems,
1.
a simple agent-based model with explicitly encoded fitness values;
2.
Avida, a self-replicating agent-based evolutionary platform (ofria2004avida); and
3.
Gen3sis, a population-level macro-ecological/evolutionary model (hagen2021gen3sis).
The following introduces each model and details experiment-specific configurations.
Simple Explicit-Fitness Model
Experiments testing the relationships between evolutionary dynamics, reconstruction error, and phylogenetic structure required a model system amenable to direct, interpretable tuning of ecology, spatial structure, and selection pressure. Finally, a parsimonious and generic model system was desired so that findings would better generalize across digital evolution systems. The core of this work relies on a simple agent-based model devised to fulfill these objectives.
Genomes in the simple explicit-fitness model comprised a single floating-point value, with higher magnitude corresponding to higher fitness. Population size 32,768 () was used for all experiments. Selection was performed using tournament selection with synchronous generations. Treatments’ selection pressure was controlled via tournament size. Mutation was applied after selection, with a value drawn from a unit Gaussian distribution added to all genomes. Evolutionary runs were ended after 262,144 () generations. Each run required around 4 hours of compute time.
Treatments incorporating spatial structure used a simple island model. In spatially structured treatments, individuals were evenly divided among 1,024 islands and only competed in selection tournaments against sympatric population members. Islands were arranged in a one-dimensional closed ring and 1% of population members migrated to a neighboring island each generation.
Treatments incorporating ecology used a simple niche model. Population slots were split evenly between niches. Organisms were arbitrarily assigned to a niche at simulation startup and were only allowed to occupy population slots assigned to that niche. Therefore, individuals exclusively participated in selection tournaments with members of their own niche. In treatments also incorporating spatial structure, an even allotment of population slots was provided for every niche on every island. Every generation, individuals swapped niches with probability (chosen so one niche swap would be expected every 1,000 generations).
For our main experiments, we defined the following “regimes” of evolutionary conditions:
plain: tournament size 2 with no niching and no islands,
weak selection: tournament size 1 with no niching and no islands,
strong selection: tournament size 4 with no niching and no islands,
spatial structure: tournament size 2 with no niching and 1,024 islands,
weak ecology: tournament size 2 with 4 niches and niche swap probability increased 100,
ecology: tournament size 2 with 4 niches, and
rich ecology: tournament size 2 with 8 niches.
In follow-up experiments testing ecological dynamics with a spatial background, we defined the following additional evolutionary “regimes:”
plain: tournament size 2 with no niching over 1,024 islands,
weak ecology: tournament size 2 with 4 niches and niche swap probability increased 100 over 1,024 islands,
ecology: tournament size 2 with 4 niches over 1,024 islands, and
rich ecology: tournament size 2 with 8 niches over 1,024 islands.
Finally, to foster generalizability of findings, all experiments were performed with two alternate “sensitivity” variables: evolutionary length in generations and mutation operator. We saved phylogenetic snapshots at 32,768 generation epochs. This allowed us to test shorter runs of 32,768 and 98,304 generations (through epochs 0 and 2) in addition to the full-length runs (through epoch 7). One additional mutation operator was tested to contrast the unit Gaussian distribution: the unit exponential distribution. Under this distribution, deleterious mutations are not possible and large-effect mutations are more likely.
Across all experiments, each treatment comprised 50 independent replicates. Testing over repeat simulations allowed us to characterize the influence of evolutionary dynamics on phylogenetic structure relative to background stochasticity, including the amount of distributional overlap in phylogeny metrics between treatments (i.e., Cliff’s delta statistic, discussed below).
Avida Model
Avida is a virtual agent-based model system used for sophisticated in silico evolution experiments (ofria2004avida). Notably, unlike the simple model described above, fitness within Avida arises implicitly from agents’ self-replication activity, in relation to other agents’ self-replication and the availability of shared exogenous resources. Within Avida, digital organisms’ genomes comprise a sequence of virtual CPU instructions. Avida conducts open-ended evaluation of each agent’s genetic program, affording the opportunity of copying itself into output memory and thereby creating an offspring agent. Replication imposes an intrinsic baseline level of copy errors (i.e., mutations), ensuring an ongoing supply of genetic variation. Because self-replicators compete for a limited quantity of population slots, Darwinian evolution ensues. This scheme induces a complex fitness landscape, which is thought to better reflect the character of biological evolution (adami2006digital).
Under baseline conditions, Avidians compete on the basis of self-replication efficiency. However, Avida may be configured to award extra CPU cycles to agents that complete boolean logic tasks. To create the potential for multiple niches, we associated each task with a depletable resource. Avidians only recieve CPU cycles for completeing a task if the corresponding resource is available, and completing a task depletes the associated resource. Importantly, tasks can each be associated with different resources, or they can all be associated with the same resource. Association of each task to an independent resource creates potential for stable ecological coexistence between task-specialized clades. We added the constraint that Avidians may harvest at most two tasks with the intention of preventing generalists from disrupting task specialization.
Our baseline “plain” treatment rewarded four boolean logic tasks (echo, not, nand, and), all drawing from a single resource with inflow rate 400 units per update. Population structure was well-mixed and, to somewhat weaken the fecundity of high-fitness Avidians, replication destroyed the parent organism with probability 0.2.
We surveyed six additional “regimes” of evolutionary conditions, defined accoring to pre-existing Avida configuration options. Notably, compared to the simple model above, these manipulations are somewhat more indirect and, therefore, potentially weaker in nature. We anticipate they might better reflect the character of variations in evolutionary drivers that might be encountered in practice.
Regimes were configured as follows,
weak selection: the probability of parent death was increased to 0.5;
strong selection: parent death probability was set to 0.0 and offspring were protected from replacement by their own offspring;
spatial structure: population was arranged as a two-dimensional toroidal grid, with Avidians replicating exclusively between neighboring population sites;
weak ecology: tasks were assigned independent resource pools, but three resources were only supplied at an inflow rate of 33 units while the fourth was supplied at 300 units per update,
ecology: each task drew from an independent resource pool with inflow rate of 100 units, and
rich ecology: 28 tasks were rewarded, each drawing from an independent resource pool with an inflow rate of 100 units.
In follow-up experiments, we defined the following additional evolutionary “regimes” with two-dimensional toroidal spatial structure,
plain: four boolean logic tasks drawing from a single resource with inflow rate 400 units per update,
weak ecology: tasks were assigned independent resource pools, but three resources were only supplied at an inflow rate of 33 units while the fourth was supplied at 300 units per update,
ecology: each task drew from an independent resource pool with inflow rate of 100 units, and
rich ecology: 28 tasks were rewarded, each drawing from an independent resource pool with an inflow rate of 100 units.
Experimental treatments using Avida comprised 30 replicates, conducted with population size 3,600 for durations of 100,000 time steps (c. 20k generations; range 9k-40k).
In order to achieve the phylogenetic tracking necessary for our experiments, we used a fork of Avida originally developed for research on MODES (dolson2019modes), linked in supplemental materials described below. This tracking system allowed two configurations: (1) tracking on the level of individual agents, where each Avidian constituted a taxonomic unit, or (2) tracking on the level of genotypes, where clonal sets of Avidians with identical genotypes constituted a taxonomic unit. Figure 1 compares example phylogenies from Avida under the “plain” treatment, tracked at organism and genome level. Surprisingly, we noticed several sign-change differences between individual-level and genotype-level tracking with respect to treatments’ effects on phylometrics (Figures 5(d) and 9). Much, but not all, of the difference related to effects of spatial structure. These differences may be in part related to occurrences of polytomies within genotype-level tracked phylogenies, as arbitrarily resolving polytomies into sets of bifurcating nodes gave somewhat more similar results to individual-level tracking. Here, we report results from individual-level tracking, which corresponds to how phylogenies were tracked in the simple model.
Gen3sis Model
In contrast to Avida and the simple model, which are agent-based, Gen3sis is a population-level model. Gen3sis abstracts evolution to interactions between spatially-dispersed, speciating subpopulations. Co-located subpopulations of different species compete for shares of site-specific carrying capacity. Interaction between subpopulation traits and environmental factors (e.g., aridity, temperature) mediates abundance determinations. Disjoined subpopulations of the same species accumulate genetic incompatibilities absent sufficient gene flow and, past a defined incompatibility threshold, speciate. Within this model, the taxonomic unit of phylogeny is species.
Gen3sis also stands out in its intended level of direct biological realism. The software can be configured to reflect the particular geological and climate histories of continental biomes. For our experiments, we used curated raster files depicting conditions over 30 million years in South America bundled with the package that had been synthesized from a number of sources (straume2020global; westerhold2020astronomically; fick2017worldclim; hagen2019mountain; annan2013new; cramwinckel2018synchronous; evans2018eocene; hollis2019deepmip; hutchinson2018climate; keating2011warm; sijp2014role; zhang2019evolution).
Among other features, Gen3sis allows differentiation of aquatic (e.g., lakes, ocean) zones from terrestrial regions with respect to suppression of inter-population migration rates. To enhance potential for manipulable spatial structure within the model, we converted 30 randomly-selected thin linear segments spanning the map to be aquatic rather than terrestrial (i.e., “rivers”). Spatial structure was then induced by adjusting the water traversal cost function.
We assessed the following evolutionary regimes,
plain: aquatic terrain imposed no additional migration penalties and population traits did not influence abundances;
spatial structure: aquatic terrain imposed a migration barrier that of terrestrial terrain;
ecology: species abundances were determined according to how close populations’ trait value matches to site temperature; and
ecology + spatial structure: the aquatic dispersal penalty was applied, and site temperature was used to determine species’ abundances.
Figure 2 shows an example Gen3sis phylogeny from the “plain” regime.
Gen3sis treatments comprised 30 replicates. Note that, unlike other surveyed models, Gen3sis does not operate with constant “population size.” Rather, species count increased continuously over each of the 30 simulated 1 million year time steps as a consequence of ongoing speciation. Maximum species count per spatial site was configured as 2,500 and within the entire simulation as 25,000, although neither limit was ever reached.
Full Gen3sis configuration files are linked in supplemental materials, described below. For more on Gen3sis itself, see (hagen2021gen3sis).
Hereditary Stratigraphic Annotations and Tree Reconstruction
Experiments testing the impact of phylogenetic inference error on phylometrics employ the recently-developed “hereditary stratigraphy” technique to facilitate phylogenetic inference (moreno2022hstrat). This technique works by attaching heritable annotations to individual digital genomes. Every generation, a new random “fingerprint” is generated and appended to the individuals’ inherited annotations. To reconstruct phylogenetic history, fingerprints from extant organisms’ annotations can be compared. Where two organisms share identical fingerprints along the record, they likely shared common ancestry. Mismatching fingerprints indicate a split in compared organisms’ ancestry. Although extensions of hereditary stratigraphy to sexual lineages are possible (moreno2024methods), all lineages used for experiments were asexual in nature.
Hereditary stratigraphy enables a tunable trade-off between annotation size and estimation accuracy. Fingerprints may be discarded to decrease annotation size at the cost of reduced density of reference points to test for common (or divergent) ancestry along organisms’ generational histories.
We test four levels of fingerprint retention. Each level is described as a “resolution” meaning that the generational distance between reference points any number of generations back is less than . So, a high percentage indicates coarse resolution and a low percentage indicates fine resolution. In detail, at the conclusion of 262,144 generation evolutionary runs,
at 33% resolution 68 fingerprints are retained per genome,
at 10% resolution 170 fingerprints are retained per genome,
at 3% resolution 435 fingerprints are retained per genome, and
at 1% resolution 1,239 fingerprints are retained per genome.
This work uses 1 byte fingerprints, which collide with probability . Greater space efficiency could be achieved using 1 bit fingerprints. However, this would require careful accounting for ubiquitous generation of identical fingerprints by chance and is left to future work.
Previous work with hereditary stratigraphy used UPGMA distance-based reconstruction techniques (moreno2022hereditary). Large-scale reconstructions required for these experiments necessitated development of a more efficient technique that did not require all pairs (i.e., ) distance comparison. To accomplish this, we devised an agglomerative tree building algorithm that works by successively adding leaf organism annotations and percolating them down from the tree root along the tree path of internal nodes consistent with their fingerprint sequence, then affixing them where common ancestry ends. This new tree-building approach reduced compute time from multiple hours to around 5 minutes in most cases. Supplementary Listing LABEL:lst:build_tree_trie provides source code with full implementation details, see (moreno2024analysis) for a more detailed discussion.
To assess the efficacy of the new agglomerative tree-building approach, we calculated all reconstructed trees’ quartet distance to their respective reference. Quartet distance ranges from 0 (between identical trees) to 0.75 (between random trees), providing in this case a measure of reconstruction error. As expected, this measure of reconstruction error varied significantly with resolution for trees across all evolutionary regimes (Kruskal-Wallis tests; all ; Supplementary Table S1). Reconstruction error also varied significantly with evolutionary regime for each reconstruction resolution level (Kruskal-Wallis tests; all ; Supplementary Table S1).
For 3% and 1% resolutions, mean reconstruction error was less than 0.01 in all cases and at 10% resolution mean reconstruction error was less than 0.05 in all cases. At 33% resolution, mean reconstruction error was less than 0.12 in all cases. The largest reconstruction errors observed at 1%, 3%, 10%, and 33% resolutions were, respectively, 0.051 (weak selection regime), 0.093 (weak 4 niche ecology regime), 0.14 (plain evolutionary regime), and 0.45 (plain evolutionary regime). Supplementary Table 27 reports mean, median, standard deviation, and maxima for reconstruction error across surveyed evolutionary conditions.
To generate reconstructed trees in experiments, we simulated the inheritance of hereditary stratigraphic annotations along a reference phylogeny to yield the set of annotations that would be attached to extant population members at the end of a run, then used our agglomerative tree building technique to infer. Thus, each reconstruction replicate has a directly-corresponding reference tree from a perfect-tree treatment replicate. Figure 3 shows a reference tree and corresponding reconstructions performed using 1%, 3%, 10%, and 33% resolution hereditary stratigraph annotations.
Phylometrics
A wide range of metrics exists for quantifying the topology of a phylogeny. Tucker et al. showed that these metrics can be classified into the following three dimensions: richness, divergence, and regularity (tuckerGuidePhylogeneticMetrics2017). Richness metrics quantify the amount of phylogenetic diversity/evolutionary history represented by a phylogeny. Divergence metrics quantify how different the units of the phylogeny are from each other. Regularity metrics quantify the variance of other properties (i.e. how consistent they are across the phylogeny). Here, we focus on four metrics spread across these categories:
Sum Pairwise Distance: This measurement sums the nodeal distance between each pair of leaf nodes. It is a metric of phylogenetic richness (tuckerGuidePhylogeneticMetrics2017), also referred to as F (izsakLinkEcologicalDiversity2000). Consequently, we would expect it to be increased by the presence of ecology or spatial structure, as both these factors increase diversity.
Colless-like Index: The original Colless Index (collessReviewPhylogeneticsTheory1982), also often referred to as (shaoTreeBalance1990), is a measure of tree imbalance (i.e., it gets higher as the tree gets less balanced). In the context of Tucker et al.’s framework, it is a regularity metric. However, the traditional Colless Index only works for strictly bifurcating trees. As our trees have multifurcations, we instead use the Colless-like Index, which is an extension of the Colless Index to multifurcating trees (mirSoundCollesslikeBalance2018). Tree imbalance is thought to be associated with varying ecological pressures (chamberlainPhylogeneticTreeShape2014; burressEcologicalOpportunityAlters) and has also been observed to increase in the presence of spatial structure (scottInferringTumorProliferative2020).
Mean Pairwise Distance: This metric is calculated by computing the shortest distance between all pairs of leaf nodes and taking the mean of these values (webbExploringPhylogeneticStructure2000). Note that these distances are measured in terms of the number of nodes in between the pair, not in terms of branch lengths. Mean pairwise distance is a metric of evolutionary divergence (tuckerGuidePhylogeneticMetrics2017). Mean pairwise distance should be increased by scenarios that promote the long-term maintenance of distinct phylogenetic branches, such as ecology. Conversely, factors that act to reduce diversity should also reduce mean pairwise distance.
Mean Evolutionary Distinctiveness: Evolutionary distinctiveness is a metric that can be calculated for individual taxa to quantify how evolutionarily different that taxon is from all other taxa in the phylogeny (isaacMammalsEDGEConservation2007). To get mean evolutionary distinctiveness, we average this value across all extant taxa in the tree. Like mean pairwise distance, mean evolutionary distinctiveness is a metric of evolutionary divergence. However, it is known to capture substantially different information than mean pairwise distance (tuckerGuidePhylogeneticMetrics2017). Unlike our other metrics, evolutionary distinctiveness is heavily influenced by branch length. We generally expect mean evolutionary distinctiveness to be increased by similar factors to mean pairwise distance.
Effect-size Analysis
We expect statistical tests for between treatments we respect to evolutionary dynamics of interest to serve as an important use case for phylometric analyses in digital evolution. As such, we wish to report the capability of reported phylometrics to discern between surveyed evolutionary conditions.
Cliff’s delta provides useful nonparametric means for such effect size analysis. This statistic reports the proportion of distributional non-overlap between two distributions, ranging from -1/1 if two distributions share no overlap to 0 if they overlap entirely (meissel2024using; cliff1993dominance). When reporting effect size, we use conventional thresholds of 0.147, 0.33, and 0.474 to distinguish between negligible, small, medium, and large effect sizes (hess2004robust).
Note that the Cliff’s delta statistic tops/bottoms out entirely once two distributions become completely separable. Although this property suits most analyses performed, it is occasionally useful to distinguish the extent of divergence between phylometric distributions past the point of complete separability. For these purposes, we perform a simple procedure to normalize phylometrics relative baseline conditions by subtracting out the baseline mean and dividing by the baseline standard deviation.
We typically pair effect-size analysis with Mann-Whitney U testing in order to assess the extent to which differences between phylometric readings under different conditions are, or are not, evidenced by available data (mann1947on). As a final detail, note that we typically report negated Cliff’s delta values where necessary to ensure positive values correspond to larger phylometric values and vice versa.
Software and Data Availability
Software, configuration files, and executable notebooks for this work are available at doi.org/10.5281/zenodo.10896667. Data and supplemental materials are available via the Open Science Framework https://osf.io/vtxwd/ (foster2017open). Note that materials for earlier versions of this work are also contained in these repositories (moreno2023toward).
All hereditary stratigraph annotation, reference phylogeny generation, and phylogenetic reconstruction tools used in this work are published in the hstrat Python package (moreno2022hstrat). This project can be visited at https://github.com/mmore500/hstrat.
This project uses data formats and tools associated with the ALife Data Standards project (lalejini2019data) and benefited from many pieces of open-source scientific software (ofria2020empirical; sand2014tqdist; 2020SciPy-NMeth; harris2020array; reback2020pandas; mckinney-proc-scipy-2010; sukumaran2010dendropy; cock2009biopython; dolson2024phylotrackpy; torchiano2016effsize; waskom2021seaborn; hunter2007matplotlib; moreno2024apc; moreno2024qspool; moreno2023teeplot; hagen2021gen3sis; ofria2004avida; torchiano2016effsize).
Results and Discussion
Phylometric Signatures of Evolutionary Dynamics
Generations Ago (approx.)
Generations Ago (approx.)
200,000
50,000
10,000
2,000
30
0
200,000
50,000
10,000
2,000
30
0
The feasibility of harnessing phylogenetic analysis to identify evolutionary dynamics hinges on the premise that these dynamics induce detectable structure within the phylogenetic record. Fortunately, as shown in Figure 4, dendrograms of phylogenetic histories from the different evolutionary conditions tested do indeed exhibit striking visual differences.
As a first step to characterizing the phylogenetic impact of spatial structure, ecology, and selection pressure, we tested whether surveyed evolutionary conditions exhibited detectable differences in a representative suite of four phylometrics: evolutionary distinctiveness, Colless-like index, mean pairwise distance, and sum pairwise distance. Figure 5 summarizes the distributions of each metric across surveyed conditions. Statistical tests confirmed that each phylometric exhibited significant variation among surveyed evolutionary conditions for both the simple model and Avida (Kruskal-Wallis tests; all ; per condition simple model, Avida; Supplementary Table 27).
To quantify the phylometric effects of surveyed evolutionary regimes, we performed nonparametric statistical comparisons against the “plain” baseline treatment. We used a measure of distributional overlap — Cliff’s delta — to assess effect sizes, binning into “negligible”, “small” (+), “medium” (++), and “large” (+++) effects based on conventional thresholds (hess2004robust). Significance at (*) was assessed through Mann-Whitney tests. Figure 5(b) shows nonparametric significance and effect size test results.
Summary of Phylometric Effects
Relative to the plain regime, all evolutionary regimes in the simple model depress the Colless-like index. Reduction in this statistic indicates that all deviations from baseline conditions increased regularity in generated phylogenies. This observation runs somewhat counter to prior results on similar tree balance metrics, in which the presence of spatial structure increased imbalance (scottInferringTumorProliferative2020). One possible contributing factor is that taxa in our phylogenies were individuals, whereas Scott et al. used genotype-level abstraction (i.e., their trees were gene trees). This possible effect of taxonomic unit is consistent with our results from the species-level phylogenies in the Gen3sis system (Figure 8), in which ecological and spatial conditions elevated Colless imbalance. Avida individual-level phylogenies were more consistent with the simple model than with Gen3sis; Colless index was significantly depressed under ecological regimes, and weakly but insignificantly depressed under spatial structure. However, other modes of evolution did not meaningfully affect Colless-like index of Avida phylogenies.
Colless-like index is sensitive to changes in evolutionary conditions. However, it appears to be the least useful metric in distinguishing different drivers of evolutionary dynamics, decreasing significantly under all non-plain evolutionary conditions.
Mean evolutionary distinctiveness was significantly higher under weak selection and with spatial structure than in the plain regime. This metric significantly decreased under strong selection and under ecological regimes, but the numerical magnitudes of these effects were relatively smaller (Figure 6). We observed similar results in Avida, except no significant effect of strong selection and weak ecology was detected on the phylometric outcome.
Mean pairwise distance was significantly depressed under all regimes except ecology and weak ecology, although again the numerical magnitude of effects on ecological regimes were relatively smaller (Figure 6). Within Avida, weak but insignificant depressing effects were observed under weak selection and spatial structure regimes. In contrast to results from the simple model, the strongest depressing effects were observed under the weak ecology and strong ecology regimes.
Finally, sum pairwise distance was significantly increased under all regimes compared to baseline, except for the strong selection regime where it was significantly depressed. Effect size was again strongest under spatial structure and weak selection (Figure 6). Avida gave similar results, except that no significant effect was detected from the weak ecology and strong selection treatments, with a weak but insignificant increase effect detected under strong selection.
Discussion of Phylometric Effects
Ecological dynamics have significant influence on the surveyed phylometrics. However, the numerical magnitudes of these effects are generally weak compared to spatial structure and selection effects (Figure 6). So, it appears careful accounting for other evolutionary dynamics (i.e., selection pressure and spatial structure) will be essential to accurate detection of ecology through phylogenetic analysis. Mean pairwise distance may play a role in identifying ecological dynamics, as ecological dynamics — in contrast to other factors such as spatial structure and changes in selection pressure — have weaker effects on this phylometric. Other phylogenetic metrics may also be better suited to detecting ecological dynamics (e.g. the ecology metric in (dolson2019modes)).
Phylometric outcomes within Avida generally mirror the simple fitness model, although in many cases phylometric effects are weaker or not significant. These less pronounced effects are not unexpected — whereas the simple model was explicitly designed for direct manipulation of evolutionary drivers, we use more subtle configurations to impose evolutionary drivers on the Avida model (particularly, with respect to selection pressure). Notably, the strong selection treatment resulted in no significant effects on phylometric values. Weak selection significantly increased mean evolutionary distinctiveness and sum pairwise distance, as with the simple model, but the effect size was small. However, spatial structure’s effect size on these metrics was large and agreed with the simple model. The ecology and rich ecology treatments agreed in sign with the simple model and generally had large effect size. However, unlike the simple model, Avida’s sole phylometric outcome under weak ecology was a strong, significant decrease in mean pairwise distance. In contrast, the simple model exhibited no effect on this metric under the weak ecology treatment.
We additionally performed a sensitivity analysis for results from the simple model over an alternate exponential mutation operator and earlier phylogeny sampling time points. The effects of evolutionary conditions on phylometrics were generally consistent with Figure 5(b) across surveyed conditions (Supplementary Figures 12 and 20).
Phylometric Signatures of Ecological Dynamics in Spatially Structured Populations
The effects of spatial structure are of particular interest, as evolution within very large populations typically entail elements of spatial structure owing to dispersal across geographic terrain. Even within in silico contexts, populations too large for a single processor will almost inevitably integrate spatial structure that reflects practical limitations of distributed computing hardware (ackley2014indefinitely; moreno2021conduit). Therefore, understanding the background effects of spatial structure on the phylogenetic signatures of other evolutionary dynamics will be essential to applications of phylogenetic inference in such applications. For this analysis, we chose to focus on ecological dynamics due to interest in how their relatively weak phylometric signatures would respond to the relatively strong influence of spatial structure (Figure 6).
Summary of Phylometric Effects with Background Spatial Structure
Figure 7 summarizes the distribution of surveyed phylometrics under the three surveyed ecological regimes and the control non-ecological regime, all with spatial population structure. Statistical tests confirmed that each phylometric exhibited significant variation among these evolutionary regimes, indicating the presence of detectable structural signatures in phylogenetic structure (Kruskal-Wallis tests; all ; per condition for simple model, Avida; Supplementary Tables 27 and 27).
As in prior experiments, all ecology treatments drove significant increases in mean evolutionary distinctiveness and sum pairwise distance. Also consistent with spatially unstructured results, rich ecology drove significant, large-effect depression of both Colless-like index and mean pairwise distance and ecology. However, in the presence of spatial structure, the ecology treatment depressed only mean pairwise distance. Without spatial structure, the ecology treatment depressed Colless-like index instead. Results under weak ecology differed notably from non-spatial baseline, with all phylometrics seeing significant, large-effect increases. Spatial-background rich ecology results from Avida agreed with the simple model. However, significant phylometric effects were not detected from the ecology and weak ecology treatments under spatial structure conditions.
For these experiments, we again performed a sensitivity analysis over an alternate exponential mutation operator and earlier phylogeny sampling time points. We found the effects of evolutionary conditions on phylometrics to be generally consistent across surveyed conditions (Supplementary Figure 19 and Supplementary Tables 27 and 27).
Species-level Phylogenies from Gen3sis Model
Significant, large-effect changes were detected across all four phylometrics in each treatment in Gen3sis (see Figure 8). However, with the exception of sum pariwise distance, effect signs of treatments were opposite to those for individual-level phylogenies from Avida and the simple model across all phylometrics. Tip count effects seem likely to play a role in this discrepancy. Unlike Avida and the simple model, which held population size constant, species richness grew freely under the Gen3sis model. It is also possible that phylometric outcomes may be sensitive to granularity level of the taxonomic unit of the phylogeny. As noted in our methods and shown in Figure 9, in Avida experiments, phylogenies using genome-level tracking (as opposed to individual-level tracking) were notably different than those using individual-level tracking. These differences also included changes in the sign of some treatment effects, lending credence to the idea that differences between Gen3sis and the individual-level phylogenies could be partially the result of Gen3sis having a more abstract taxonomic unit.
Phylometric Bias of Reconstruction Error
Shifting from perfect phylogenetic tracking to approximate phylogenetic reconstruction will facilitate efficiency and robust digital evolution simulations at scale, but introduces a complicating factor into phylogenetic analyses: tree reconstruction error. A clear understanding of the impact of these errors on the computed phylometrics will be necessary for informative future phylogenetic analyses.
To explore this question, we compared phylometrics computed on reconstructed trees to corresponding true reference trees under the simple model (Wilcoxon tests; per condition; Supplementary Table 27). To err towards conservatism in detecting phylometric biases, we did not correct for multiple comparisons. Reconstructions were performed across a range of precisions, ranging from 1% relative resolution for MRCA estimates (most precise) to 33% relative resolution for MRCA estimates (least precise). Precision was manipulated by adjusting the information content of underlying hereditary stratigraphic genome annotations used to perform phylogenetic reconstruction (moreno2022hereditary). Note that important differences exist the between nature of reconstruction error under hereditary stratigraphy versus traditional biosequence-based methods, discussed further below.
Phylometric Sensitivites to Reconstruction Error
For each phylometric, we sought to determine the minimum resolution required to achieve statistical non-detection (i.e., ) of bias between reconstructions and their corresponding references. For nearly all cases, 3% reconstruction resolution was sufficient to achieve statistical indistinguishability between reference and reconstruction. Mean evolutionary distinctiveness and sum pairwise distance were particularly robust to reconstruction error, showing no detectable bias even at only 33% reconstruction resolution.
Phylometric sensitivity to reconstruction error was broadly consistent across evolutionary regimes. Figure 10 summarizes these results.
Where detectable, estimation uncertainty bias decreased all surveyed phylometrics’ numerical value. So, when testing for expected increases in phylometric values, the potential for systematic false positives due to reconstruction error can be discounted. Supplementary Figure 23 provides a full comparison of the distribution of phylometric estimates on reference trees with the distributions of phylometric estimates for reconstructed trees across reconstruction resolutions.
The relationship between reconstruction error and phylometric bias was similar under spatially structured regimes, the alternate exponential mutation operator, at earlier phylogeny sampling time points, and in the Avida model system (Supplementary Figures 21, 25, 24, 26 and 27; Supplementary Table 27 and 27). Notably, however, reconstruction bias persisted at even 1% relative resolution in some conditions of the sensitivity analysis and Avida experiments. Forthcoming work has found that the byte-differentia hereditary stratigraphy configuration used for this experiment tends to lump closely contemporaneous lineage splitting events into a polytomy rather than a double-branching event — i.e., indicating uncertainty rather than introducing error. In contrast, reconstructions on single-bit differentia contain erroneously- sequenced branching rather than artifactual polytomies. So, future work should explore whether working with single bit differentia could lessen phylometric bias.
Detection of Evolutionary Drivers’ Signatures in Reconstructed Phylogenies
Our last objective was to assess how reconstruction error might affect detection of phylometric signatures induced by treatment conditions. That is, we sought to perform a sort of “integration test” for detection of treatment conditions when working with imperfect reconstructions rather than perfect phylogenies. To this end, we compared the phylometric outcomes of strong/weak selection, spatial structure, ecology, and weak/rich ecology relative to plain conditions for phylogenies reconstructed at 1%, 3%, 10%, and 33% resolution levels. Supplemental Figures 30 and 31 show heatmaps with sign, effect size, and significance of phylometric effects across gradations of reconstruction precision for the simple model and Avida, respectively. In most cases, 3% resolution sufficed to fully recover phylometric effects of treatments observed with perfectly-tracked phylogenies.
Conclusion
Because phylogenies are an abstraction that generalizes across all evolutionary processes, phylogenetic analysis can be applied across a breadth of biological and artificial life systems. Consequently, a broad set of cross-disciplinary use cases exist for inferring the processes that shaped a phylogeny by quantifying its topology. Indeed, extraction of information about evolutionary dynamics from phylogenetic history has been a longstanding and productive theme in evolutionary biology (pagel1997inferring). This work seeks to contribute in that vein, by establishing foundations necessary to assess three fundamental evolutionary drivers — spatial structure, ecology, and selection pressure — in complex, distributed evolving populations.
First, we investigated the strength and character of structural signatures in phylogenetic histories left by selection pressure, ecology, and spatial population structure. These drivers induced readily detectable effects across our suite of four surveyed phylometrics. These effects were generally consistent across surveyed individual-level phylogenies, but differed notably under genotype-level tracking and under species-level simulation. Although the directions of phylometric effects were generally consistent across treatments expected to increase phylogenetic richness (i.e., weak selection, ecologies, and spatial structure), effect significance/size for particular phylometrics differed among some treatments. Compared to results from the simple model, which was amenable to strong treatment manipulations, the more sophisticated Avida model system expressed fewer phylometric effects between treatments.
Compared to spatial structure and selection pressure, drivers designed to induce ecological coexistence exerted relatively muted, sparse effects on phylogenetic metrics. However, follow-up experiments with both individual-level model systems revealed that phylometric signatures can remain detectable under background conditions of spatial population structure. Interactions, though, are sometimes surprising. A notable example is the reduced mean pairwise distance observed in Avida under conditions combining spatial structure and ecology, but not in the baseline ecology treatment. These results highlight the complexity of how interacting evolutionary dynamics impact phylogenetic structure. Even a single dynamic in isolation can influence phylometrics in opposite directions. For instance, we found that changes to selection pressure in either direction significantly reduced trees’ Colless-like index under the simple model.
Comparison of phylometrics from perfect-fidelity trees against corresponding reconstructions revealed that phylometric statistics differed in sensitivity to reconstruction error introduced by hereditary stratigraphy-based tracking. Colless-like index and mean pairwise distance were more sensitive, and in some conditions, reconstruction persisted at even 1% reconstruction resolution. On the other hand, mean evolutionary distinctiveness and sum pairwise distance were particularly robust to reconstruction error, with no bias detected even at 33% resolution reconstruction. Results suggest 3% reconstruction resolution as a reasonable ballpark parameterization for applications of hereditary stratigraphy with downstream phylogenetic analysis. Where it did occur, the sign effect of phylometric bias was consistent, which might simplify considerations to account for it in experiments.
Our findings identify key next steps necessary for development of rigorous phylogenetic assays to test for key evolutionary drivers. Because treatments did not induce widespread opposite-sign effects, our results do not indicate a straightforward, qualitative approach to disambiguate drivers’ signatures. Future work should survey comprehensive panels of phylometrics, which might provide a stronger foothold to this end (tuckerGuidePhylogeneticMetrics2017). Alternately, more sophisticated quantitative approaches leveraging differential effect size ratios may be fruitful. Possibilities include application of machine learning techniques over broad phylometric panels (voznica2022deep) or even inference directly over phylogenetic structure itself via graph neural networks and related methods (lajaaiti2023comparison). The capability of artificial life approaches to generate high-quality data sets with high replication counts will be a significant asset to such work. More information could be harnessed by incorporating trait data (nozoe2017inferring) or by using time-sampled phylogenies (volz2013viral), which would include information about “ghost” lineages that have extincted at earlier time points.
Additionally, robust normalization methods will be crucial to enable meaningful comparisons between phylogenies differing in (1) population size, (2) population subsampling level, (3) depth of evolutionary history, and (4) demographic characteristics of organism life histories. Despite some substantial work on this front (shao1990tree; mir2018sound), significant challenges remain. It is worth noting that most analysis reported here normalizes to phylometric values sampled from “plain” evolutionary regimes, thereby implicitly depending on the availability of such data. However, stripped-down baseline condition experiments may not always be possible, especially in biological model systems. Extensive, likely sophisticated, considerations would be necessary to establish a priori phylometric value predictions that could be compared against instead of testing against empirical values. Finally, as evidenced in our experiments, reconstruction error can systematically bias phylometric values. In scenarios where reconstruction error is unavoidable, additional normalizations may be necessary. However, it is promising that some phylometrics appear thoroughly robust to reconstruction error and, in other cases, sufficient reconstruction accuracy might be achieved to obviate error bias.
The overriding objective of work presented here is to equip biologists and computational researchers with capability to investigate fundamental evolutionary drivers within large-scale evolving populations. To this end, methodology must advance hand-in-hand with the software infrastructure required to put it into practice. Artificial life experiments, specifically, tend to require particular instrumentation to collect phylogenetic histories, and we have been active in making general-use plug-and-play solutions available freely through the Python Packaging Index (moreno2022hstrat; dolson2024phylotrackpy). As each field develops, artificial life will benefit greatly from domain-specific software and methods developed within traditional evolutionary bioinformatics and, we hope, vice versa.
Acknowledgment
This research was supported in part by NSF grants DEB-1655715 and DBI-0939454 as well as by Michigan State University through the computational resources provided by the Institute for Cyber-Enabled Research. This material is based upon work supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE-1424871. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. This material is based upon work supported by the Eric and Wendy Schmidt AI in Science Postdoctoral Fellowship, a Schmidt Futures program.
S1 Supplemental Material
myTableStyle
longtable=c c c c c c c c,
before reading=,
table head= Tree Distance Metric Statistic Reconstruction Fidelity Epoch Mutation Distribution
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
head to column names,
respect all,
[…]
myTableStyle
longtable=c c c c c c c c,
before reading=,
table head= Tree Distance Metric Statistic Evolutionary Regime Epoch Mutation Distribution
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
head to column names,
respect all,
[…]
Generations Ago (approx.)
Generations Ago (approx.)
200,000
150,000
100,000
50,000
0
200,000
150,000
100,000
50,000
0
myTableStyle
longtable=c c c c c c c c,
before reading=,
table head=evolutionary regime reconstruction quality sensitivity analysis conditions quartet dist mean quartet dist median quartet dist std quartet dist max
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
head to column names,
respect all,
[…]
myTableStyle
longtable=c c c c c c c c,
before reading=,
table head=evolutionary regime reconstruction quality sensitivity analysis conditions triplet dist mean triplet dist median triplet dist std triplet dist max
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
head to column names,
respect all,
[…]
myTableStyle
longtable=c c c c c c c c,
before reading=,
table head=evolutionary regime 1 evolutionary regime 2 Phylometric Statistic Epoch Mutation Distribution
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
head to column names,
respect all,
[…]
myTableStyle
longtable=c c c c c c c c,
before reading=,
table head=evolutionary regime 1 evolutionary regime 2 Phylometric Statistic Epoch Mutation Distribution
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
head to column names,
respect all,
[…]
myTableStyle
longtable=c c c c c c c c,
before reading=,
table head=Reconstruction Fidelity Phylometric Statistic Evolutionary Regime Epoch Mutation Distribution
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
head to column names,
respect all,
[…]
myTableStyle
longtable=c c c c c c c c,
before reading=,
table head=Reconstruction Fidelity Phylometric Statistic Evolutionary Regime Epoch Mutation Distribution
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
head to column names,
respect all,
[…]
myTableStyle
longtable=c c c c c c c c,
before reading=,
table head= Phylometric Statistic Epoch Mutation Distribution
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
after reading=
head to column names,
respect all,
[…]
myTableStyle
longtable=c c c c c c c c,
before reading=,
table head= Phylometric Statistic Epoch Mutation Distribution
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
after reading=
head to column names,
respect all,
[…]
myTableStyle
longtable=c c c c c c c c,
before reading=,
table head= Phylometric Statistic Epoch Mutation Distribution
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
head to column names,
respect all,
[…]
myTableStyle
longtable=c c c c c c c c,
before reading=,
table head= Phylometric Statistic Epoch Mutation Distribution
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
Kruskal-Wallis one-way analysis of variance tests for nonequivalence of reconstruction error distributions among surveyed evolutionary regimes.
Comparisons were performed independently within each sensitivity analysis condition and reconstruction fidelities.
Error reported as both quartet and triplet distance between reconstructed tree and corresponding reference tree.
Sample size for each population, with populations (i.e., evolutionary regimes) compared within each sensitivity analysis condition.
[…]
head to column names,
respect all,
[…]
S1.1 Trie-based Phylogenetic Tree Reconstruction Implementation
This section provides source code and docstrings for trie-based phylogeny reconstruction.
The main reconstruction logic takes place in build_trie_from_artifacts (Listing LABEL:lst:build_tree_from_artifacts). First, hereditary stratigraphic annotations associated with each extant population member are sorted by the generational depth of their lineages in ascending order. This ensures that higher-per-generation-density hereditary stratigraphic records of annotations from shorter lineages are encountered first. (In this work, because generations are synchronous, this step had no effect.)
Next, annotations are inserted into a trie data structure one at a time. All annotations begin at the root node of the trie, representing a universal common ancestor. Each subsequent node in the trie denotes a particular “fingerprint” value associated with its generation of creation. (This is referred to in the source code as an “allele”). Paths traced from the root node towards the tips therefore represent a particular sequence of allele generation events; organisms will share common allele generation histories for the over the evolutionary interval they share common ancestry.
Each annotation descends the trie along the path corresponding to the sequence of alleles contained within its own record. The GetDeepestCongruousAllele member function of the TrieInnerNode class implements this process. When no further consistent alleles are present in the trie, the remaining alleles within the annotation are unrolled to create a new unifurcating branch. The label corresponding to the annotation is then appended as the leaf node on that branch. The InsertTaxon member function of the TrieInnerNode class implements this process. The TrieInnerNode and TrieLeafNode classes are provided in Listings LABEL:lst:TrieInnerNode and LABEL:lst:TrieLeafNode.
The build_tree_trie method (Listing LABEL:lst:build_tree_trie) serves as the API entrypoint for the overall tree reconstruction routine. Default options were used for all reconstructions in this work. The build_tree_trie entrypoint delegates to build_tree_trie_ensemble (Listing LABEL:lst:build_tree_trie_ensemble), which ultimately delegates to build_build_trie_from_artifacts. These additional source files are included due to docstring parameter descriptions referenced in other files and information about minor trie postprocessing performed to perform ancestor origin time estimates and convert the trie structure into a standard phylogenetic output format. Listing LABEL:lst:AssignOriginTimeNaiveTriePostprocessor provides the particular postprocessing implementation used to estimate ancestor origin times in reconstructions performed as part of this work.
The full source context for the files included in listings can be found at https://github.com/mmore500/hstrat.